OpenStudy (anonymous):

So Integral of ln x/x What method should I use? So far I used u'v+u'v and integrating by parts but it didnt work.

OpenStudy (anonymous):

the answer is 1/2 (ln (x))^2 + c

OpenStudy (anonymous):

okey I let u = 1/x probably why, okey Ill do it agian.

OpenStudy (anonymous):

oh okey, what I forgot was to integrate again after doing u'v+u'v, thanx for the quick reply!

OpenStudy (anonymous):

oh I didnt know that.

OpenStudy (anonymous):

I did it here, let u = t^2 --> du = 2t dt dv = e^t dt--> v = e^t the whole thing is \[\int_0^2 t^2e^tdt=t^2e^t( from(0 to 2)) -\int_0^2 2te^tdt\] =4e^2-\(2\int_0^2 te^tdt\) the second part only: \[2\int_0^2te^tdt\]

OpenStudy (anonymous):

let u =t --> du =dt dv=e^tdt--> v= e^t \[2\int_0^2te^tdt=2[te^t(from 0to2)-\int_0^2e^tdt\] =2[2e^2-e^2]=2e^2 replace to the whole thing above =\(4e^2-2e^2=2e^2\)

OpenStudy (anonymous):

@Xlegalize